# Navigating the Social Welfare Frontier: Portfolios for Multi-objective Reinforcement Learning

### v1

@@ 4.2 Heuristic under a Budget Constraint
$\mathcal{Q}(\Pi^{\prime}_{t})={\min}_{p\in[-\infty,1]}\frac{\max_{\pi\in\Pi^{\prime}_{t}}v^{(\ell)}(\pi,p)}{v^{(\ell)}_{*}(p)}.$ (4)
@@ 5 Numerical Experiments
In this section, we present the results of our numerical experiments. Additional details of the experimental setups and the environments are provided in Appendix B.
@@ 5.2 Environments
Taxi Environment. We consider a synthetic setting based on the works of [10, 12]. The taxi environment consists of a grid world where a taxi driver serves passengers across $N=4$ different source-destination pairs. When the agent drops off a passenger at their destination, it earns a reward corresponding to that route. However, since the taxi can only carry one passenger at a time, it must decide which route to prioritize. A fair agent should serve all source-destination pairs equitably, avoiding the neglect of more challenging routes. We train a $p$-mean maximizing policy using Welfare Q-Learning [12]. Finally, we experiment on $v^{(1)}$ and define $\Pi$ as the finite set of pre-trained policies.
@@
Healthcare Intervention. We consider a real-world healthcare intervention problem, modeled as an RMAB problem described in Section 1.2 [34]. ARMMAN [3], an NGO based in India, runs large-scale maternal and child mobile health programs. One of their initiatives delivers critical health information via weekly automated voice messages. To enhance engagement, a limited number of beneficiaries receive direct calls from health workers each week, with call assignments determined by a policy. The problem involves multiple reward functions, each prioritizing different socio-demographic groups among the beneficiaries. We conduct experiments using $v^{(2)}$, where $\Pi$ consists of a finite set of pre-trained policies. All experiments are strictly secondary analyses and adhere to ethics board approvals; for further discussion, refer to our impact statement.
@@ 5.3 Results
Size. The portfolios computed using Algorithm $p$-MeanPortfolio remain small while achieving a very high approximation factor. Across all three environments, the portfolio size never exceeds 10, yet still attains an approximation factor close to 1. This suggests that a small portfolio is sufficient to effectively cover the entire spectrum of $p$ values in $[-\infty,1]$.
@@
Table 2: A comparison of the number of oracle calls and actual approximation ratios for $p$-MeanPortfolio and BudgetConstrainedPortfolio across different portfolio sizes. While $p$-MeanPortfolio often achieves slightly better approximation, it requires significantly more oracle calls. [^3] MIT
@@
Diversity. As we have shown in Figure 1 in Section 1.2, portfolios can simplify policy deployment decisions by presenting a small yet diverse set of policies. Figure 2 in Appendix B.2 further shows the diversity of policies in the natural disaster setting, showing how different clusters are prioritized by each policy. Importantly, these are not arbitrary policies but optimal policies for specific $p$ values, with approximation guarantees extending to other $p$s as well. The portfolios effectively summarize the trade-offs within the policy space, providing a structured way to navigate decision-making.
33footnotetext: Here, we observe a drop in the actual approximation factor of $p$-MeanPortfolio when the portfolio size is 8. Although this portfolio is computed using $\alpha=0.9$, its achieved approximation is slightly lower. This discrepancy is likely due to the inherent randomness in the RL algorithm and the challenges of computing exact optimal policies in RL settings, as discussed in Section 4.
@@ Appendix A Omitted Proofs
$f(\mathbf{x},-\infty)\geq\alpha\cdot f(\mathbf{x},p)\quad\forall\ p\in[-\infty,p_{0}].$
@@
Suppose $0<x_{1}\leq\ldots\leq x_{n}$, so that $f(\mathbf{x},-\infty)=\min_{j\in[N]}x_{j}=x_{1}$. Given $p\in[-\infty,p_{0}]$, denote $q=-p\geq\frac{\ln N}{\ln(1/\alpha)}$. Then, since
@@
Given an MDP $\mathcal{M}$ with $N$ reward functions, set $\Pi$ of feasible policies, and an approximation factor $\alpha\in(0,1)$, denote $p_{0}=-\frac{\ln N}{\ln(1/\alpha)}$. Then, for a given aggregation rule $\ell\in\{1,2\}$, policy $\pi_{0}={\arg\max}_{\pi\in\Pi}v^{(\ell)}(\pi,p_{0})$ is an $\alpha$-approximation for all $p\in[-\infty,p_{0}]$. That is,
$v^{(\ell)}(\pi_{0},p)\geq v_{*}^{(\ell)}(p)=\max_{\pi\in\Pi}v^{(\ell)}(\pi,p)\quad\forall\ p\in[-\infty,p_{0}].$
@@
$=\alpha\cdot\max_{\pi\in\Pi}\cdot v^{(\ell)}(\pi,p_{0})$
@@
    For a given aggregation rule $\ell\in\{1,2\}$, define $w(p):=\ln v^{(\ell)}_{*}(p)$. Then, for all distinct $p,q\in[-\infty,1]$,
@@
Part 1. Note that $g(\beta\mathbf{x},p)=\ln\beta+g(\mathbf{x},p)$ for all $\mathbf{x}\in\mathbb{R}^{N}_{>0}$, $p\in[-\infty,1]$, and $\beta>0$. Therefore, we can assume without loss of generality that $LH=1$ and $UH=\frac{UH}{LH}=\kappa$.
@@
$Np\ \frac{dg(\mathbf{x},p)}{dp}\leq\sum_{i}\ln x_{i}\left(\frac{x_{i}^{p}}{\mu_{p}}-1\right)\leq\ln\kappa\sum_{i:x_{i}^{p}\geq\mu_{p}}\left(\frac{\mathbf{x}_{N}^{p}}{\mu_{p}}-1\right)\leq N\ln\kappa\left(\frac{\mathbf{x}_{N}^{p}}{\mu_{p}}-1\right).$ (7)
@@
Part 3. Note that $w(p)=\ln v^{(\ell)}_{*}(p)=\ln\max_{\pi\in\Pi}v^{(\ell)}(\pi,p)=\max_{\pi\in\Pi}\ln v^{(\ell)}(\pi,p)$. Given distinct $p,q\in[-\infty,1]$, denote $\pi_{p}={\arg\max}_{\pi\in\Pi}\ln v^{(\ell)}(\pi,p)$. Then, by Part 1,
@@
Lemma A.4 shows that $\pi_{0}$ is an $\alpha$-approximation for all $p\in[-\infty,p_{0}]$. It is therefore sufficient to prove that for all $t\in[0,K-1]$, policy $\pi_{t}$ is an $\alpha$-approximation for all $p\in[p_{t},p_{t+1}]$. Since $p_{t+1}=\textsc{LineSearch}(v^{(\ell)},p_{t},\Pi,\alpha)$, Lemma A.7 implies this.
@@
Appendix B Experimental Details
@@ B.1 Taxi Environment
This environment consists of a taxi agent whose task is to deliver passengers from source to destination. The world consists of a 6x6 grid and based on the environment setup of [12], we consider 4 source-destination pairs. Whenever a taxi moves to a source point, it can pick up a passenger, move to the destination and drop off the passenger, thus receiving a reward. Since some routes could be easier to serve, the agent has to decide which route to serve more often. Rewards are multi-dimensional, each dimension corresponding to each route. We consider the following source-destination pairs: $source\_coords=[[0,0],[0,5],[3,0],[1,0]]$, $destination\_coords=[[1,5],[5,0],[3,3],[0,3]]$
@@
The state space consists of information on the location of the taxi, the locations of passengers, and whether passengers have been picked up by the taxi at the moment.
@@
The agent can take one of 6 actions: move north, south, east, west, and pick up or drop off a passenger in the current grid cell.
@@
The reward function gives reward 0 for moving, -10 reward for taking an invalid action of picking up or dropping off a passenger at the wrong coordinates, and 30 reward for dropping a passenger at the right destination.
@@
Given a scalarization function (or welfare function ), the task is to find a policy $\pi^{*}$ that maximizes the expected value of the scalarized return. Specifically:
@@
We learn this policy using the Welfare Q-Learning algorithm proposed in [12]. For every problem setting, we train the policy for 200 episodes. Every episode is a finite horizon problem with 1000 timesteps. We consider the discount factor $\gamma=0.99$.
B.2 Natural Disaster
@@ 12 Low Near Middle-Income 230 100
Portfolio of Policies
We showcase one of the portfolios for this setting obtained by Algorithm 1 in Figure 2.
B.3 Healthcare Intervention
The Healthcare Intervention problem is based on the large-scale mobile-health program run by the NGO [3]. The goal of the program is to maximize the engagement of beneficiaries with the program using limited service call interventions. Based on previous works, we model this problem as a Restless Multi-Armed Bandit (RMAB) problem where we have N arms or beneficiaries and a budget of K service calls to give every week. For every beneficiary, we have information on their listenership with the voice call. This is considered as an engaging or non-engaging state if the listenership is above or below 30 seconds threshold respectively. One week of time is considered as one timestep. Finally, every week, the action can be to place (active) or not to place a live service call (passive). Additionally, for every beneficiary, we have information on their socio-demographic characteristics such as age, income, and education. We use real-world data from service quality improvement conducted by ARMMAN in January 2022 [33]. Further, for our experiments, we sample data for $N=2100$ beneficiaries and run the experiment for $H=12$ timesteps.
Reward Design
The default reward incentivizes agents to be in engaging state. Thus, agents receive reward 1 if they are in an engaging state, and 0 otherwise. However, due to varying policy needs over time or over different geographies, health workers often have to prioritize some beneficiaries more than others. To design reward functions based on natural language preferences, we use the Social Choice Language Model [32], which proposes a set of reward functions aligned with the preference. Finally, it outputs the learned policies corresponding to every reward function. We used 200 reward functions ($|\Pi|=200$).

### v2

@@ 4.2 Heuristic under a Budget Constraint
$\mathcal{Q}(\Pi^{\prime}_{t})={\min}_{p\leq 1}\frac{\max_{\pi\in\Pi^{\prime}_{t}}v^{(\ell)}(\pi,p)}{v^{(\ell)}_{*}(p)}.$ (4)
@@ 5 Numerical Experiments
In this section, we present the results of our numerical experiments. Additional details of the experimental setups and the environments are provided in Appendix E. The code for our experiments can be found at https://github.com/jaimoondra/approximation-portfolios-for-rl/.
@@ 5.2 Environments
Taxi Environment. We consider a synthetic setting based on the works of Dietterich (2000); Fan et al. (2022). The taxi environment consists of a grid world where a taxi driver serves passengers across $N=4$ different source-destination pairs. When the agent drops off a passenger at their destination, it earns a reward corresponding to that route. However, since the taxi can only carry one passenger at a time, it must decide which route to prioritize. A fair agent should serve all source-destination pairs equitably, avoiding the neglect of more challenging routes. We train a $p$-mean maximizing policy using Welfare Q-Learning (Fan et al., 2022). Finally, we experiment on $v^{(1)}$ and define $\Pi$ as the set of all possible policies in the MDP.
@@
Healthcare Intervention. We consider a real-world healthcare intervention problem, modeled as an RMAB problem described in Section 1.2 (Verma et al., 2024b). ARMMAN (ARMMAN, 2024), an NGO based in India, runs large-scale maternal and child mobile health programs. One of their initiatives delivers critical health information via weekly automated voice messages. To enhance engagement, a limited number of beneficiaries receive direct calls from health workers each week, with call assignments determined by a policy. The problem involves $N=59$ reward functions, each prioritizing different socio-demographic groups among the beneficiaries. We conduct experiments using $v^{(2)}$, where $\Pi$ consists of a finite set of pre-trained policies. All experiments are strictly secondary analyses and adhere to ethics board approvals; for further discussion, refer to our impact statement.
@@ 5.3 Results
Size. The portfolios computed using Algorithm $p$-MeanPortfolio remain small while achieving a very high approximation factor. Across all three environments, the portfolio size never exceeds 10, yet still attains an approximation factor close to 1. This suggests that a small portfolio is sufficient to effectively cover the entire spectrum of $p$ values $\leq 1$.
@@
However, BudgetConstrainedPortfolio and random $p$ sampling occasionally outperform $p$-MeanPortfolio. This is because $p$-MeanPortfolio selects policies solely to meet the input approximation factor $\alpha$, but has no incentive or mechanism to surpass this approximation quality.
In contrast, BudgetConstrainedPortfolio fully utilizes the input budget $K$, and strategically selects boundary points first (a small initial $p$ followed by $p=1$). When $K$ is very small (1 or 2), this initial heuristic strategy can provide more effective coverage across $p\leq 1$, explaining its occasional superior performance. Likewise, in rare cases where $K=1$, a randomly chosen $p$ may happen to yield better coverage than the single policy selected by $p$-MeanPortfolio.
@@
Table 2: A comparison of the number of oracle calls and actual approximation ratios for $p$-MeanPortfolio and BudgetConstrainedPortfolio across different portfolio sizes. While $p$-MeanPortfolio often achieves slightly better approximation, it requires significantly more oracle calls.  [^3] Here, we observe a drop in the actual approximation factor of $p$-MeanPortfolio when the portfolio size is 8. Although this portfolio is computed using $\alpha=0.9$, its achieved approximation is slightly lower. This discrepancy is likely due to the inherent randomness in the RL algorithm and the challenges of computing exact optimal policies in RL settings, as discussed in Section 4.
@@
Diversity. As we have shown in Figure 1 in Section 1.2, portfolios can simplify policy deployment decisions by presenting a small yet diverse set of policies. We also provide the entire $p$ values chosen by $p$-MeanPortfolio in Appendix D, Table 4. We observe that the $p$ values are quite diverse, which corroborates the algorithm’s objective to cover the entire $p$ range only with these selected values. Figures 3 and 4 further illustrate how the outcomes of the optimal policies for different $p$ values in the portfolio lead to varying impacts for the stakeholders. Importantly, these are not arbitrary policies but optimal policies for specific $p$ values, with approximation guarantees extending to other $p$s as well.
@@ 6 Conclusion
Acknowledgments
We thank NSF AI Institute for Societal Decision Making (AI-SDM) Award No. 2229881, ONR MURI N00014-24-1-2742, and NSF CAREER 2239824 for their support.
Impact Statement
Our work studies the setting where a deployed RL policy impacts multiple stakeholders differently, a scenario that arises in various societal applications, including healthcare and LLMs. By summarizing the space of policies induced under different $p$-values, our proposed approach aims to assist decision-makers in making informed and equitable choices. Furthermore, this work has the potential to stimulate discussions within the machine learning community about the complexities and pluralities of social welfare notions, and the importance of addressing them. However, our focus is limited to $p$-means social welfare functions, which may not fully capture all dimensions of fairness in real-world scenarios.
Our experiments use real-world healthcare data provided by the NGO ARMMAN. These experiments are strictly secondary analyses, with no real-world deployment of the proposed algorithm or baseline methods. All data usage complied with ARMMAN’s ethics board approvals, and data exchange and analysis followed their ethics review committee’s guidelines, including privacy, informed consent, and anonymization. All of the analyses is conducted in close collaboration with ARMMAN.
This work belongs to the broader area of algorithmic decision-making, where deployment decisions can have real-world implications. We emphasize the importance of not relying on this study in isolation. Responsible deployment requires comprehensive field testing and human oversight. Moreover, we acknowledge broader challenges inherent in algorithmic decision-making, including potential biases in data, inaccuracies in the underlying mathematical models, accountability issues, and concerns related to consent and privacy. Addressing these challenges is essential to ensuring that algorithmic decisions are ethically sound.
@@ Appendix A Omitted Proofs
$f(\mathbf{x},-\infty)\geq\alpha\cdot f(\mathbf{x},p)\quad\forall\ p\leq p_{0}.$
@@
Suppose $0<x_{1}\leq\ldots\leq x_{n}$, so that $f(\mathbf{x},-\infty)=\min_{j\in[N]}x_{j}=x_{1}$. Given $p\leq p_{0}$, denote $q=-p\geq\frac{\ln N}{\ln(1/\alpha)}$. Then, since
@@
Given an MDP $\mathcal{M}$ with $N$ reward functions, set $\Pi$ of feasible policies, and an approximation factor $\alpha\in(0,1)$, denote $p_{0}=-\frac{\ln N}{\ln(1/\alpha)}$. Then, for a given aggregation rule $\ell\in\{1,2\}$, policy $\pi_{0}={\arg\max}_{\pi\in\Pi}v^{(\ell)}(\pi,p_{0})$ is an $\alpha$-approximation for all $p\leq p_{0}$. That is,
$v^{(\ell)}(\pi_{0},p)\geq v_{*}^{(\ell)}(p)=\max_{\pi\in\Pi}v^{(\ell)}(\pi,p)\quad\forall\ p\leq p_{0}.$
@@
$=\alpha\cdot\max_{\pi\in\Pi}\cdot v^{(1)}(\pi,p_{0})$
@@
    For a given aggregation rule $\ell\in\{1,2\}$, define $w(p):=\ln v^{(\ell)}_{*}(p)$. Then, for all distinct $p,q\leq 1$,
@@
Part 1. Note that $g(\beta\mathbf{x},p)=\ln\beta+g(\mathbf{x},p)$ for all $\mathbf{x}\in\mathbb{R}^{N}_{>0}$, $p\leq 1$, and $\beta>0$. Therefore, we can assume without loss of generality that $LH=1$ and $UH=\frac{UH}{LH}=\kappa$.
@@
$Np\ \frac{dg(\mathbf{x},p)}{dp}\leq\sum_{i}\ln x_{i}\left(\frac{x_{i}^{p}}{\mu_{p}}-1\right)\leq\ln\kappa\sum_{i:x_{i}^{p}\geq\mu_{p}}\left(\frac{x_{N}^{p}}{\mu_{p}}-1\right)\leq N\ln\kappa\left(\frac{x_{N}^{p}}{\mu_{p}}-1\right).$ (7)
@@
Part 3. Note that $w(p)=\ln v^{(\ell)}_{*}(p)=\ln\max_{\pi\in\Pi}v^{(\ell)}(\pi,p)=\max_{\pi\in\Pi}\ln v^{(\ell)}(\pi,p)$. Given distinct $p,q\leq 1$, denote $\pi_{p}={\arg\max}_{\pi\in\Pi}\ln v^{(\ell)}(\pi,p)$. Then, by Part 1,
@@
We note that this size bound can be tightened slightly to $O\left(\frac{(\ln\kappa)(\ln\kappa+\ln\ln N)}{\ln(1/\alpha)\ln\ln(1/\alpha)}\right)$; we present the the above cleaner bound for simplicity.
@@
Lemma A.4 shows that $\pi_{0}$ is an $\alpha$-approximation for all $p\leq p_{0}$. It is therefore sufficient to prove that for all $t\in[0,K-1]$, policy $\pi_{t}$ is an $\alpha$-approximation for all $p\in[p_{t},p_{t+1}]$. Since $p_{t+1}=\textsc{LineSearch}(v^{(\ell)},p_{t},\Pi,\alpha)$, Lemma A.7 implies this.
@@
Appendix B Generalized $p-$Means: Illustrative Example
To clarify the meaning behind $p$-means, we constructed an illustrative example in Figure 2. In this example, three vectors $A$, $B$ and $C$ represent different policies, where each vector’s $i$th entry is the total reward for the $i$th stakeholder. Here, $A$ is a balanced policy, $B$ is unbalanced with the largest sum, and $C$ lies in between. Figure 2 (c) plots the $p$-mean values of these vectors, showing that varying $p$ leads to different optimal policy.
(a) Entries of $A,B,C$
(b) $p-$means at selected $p$ values.
(c) $p-$means across different $p$ values.
Figure 2: Side-by-side comparison of entries and $p$-means for vectors $A,B,C$.
Appendix C Comparison with other MORL Algorithms
[Yu et al., 2024] optimize a generalized Gini welfare assuming a fixed weight vector. We address a fundamentally different scenario, where decision-makers are uncertain about the appropriate fairness criterion (e.g., the choice of $p$ in $p$-means or the selection of weights in Gini welfare) in advance. Optimizing a single policy for one fairness parameter can lead to poor outcomes under different criteria (see the approximation qualities of random baselines in our experiments). Our method computes a small, representative set of policies covering the entire spectrum of fairness criteria. This allows decision-makers to select confidently from this set, without worrying about suboptimality under other potential choices of $p$.
[Yang et al., 2019] focuses on weighted linear objectives and learns a single policy, whereas we focus on $p$-means and compute a portfolio (multiple policies). [Alegre et al., 2023] builds portfolios focusing on weighted linear objectives (as opposed to $p-$means) and does not offer guarantees on portfolio size or oracle complexity. [Reymond et al., 2022] trains a single policy to approximate the Pareto frontier, which does not directly correspond to any social welfare function. To the best of our knowledge, our work is the first to (1) propose a multi-policy approach in $p$-means and (2) provide theoretical guarantees on portfolio size, approximation quality, and oracle complexity simultaneously.
In Table 3, we also present comparisons between $p$-MeanPortfolio and the Generalized Policy Improvement (GPI) algorithm of [Alegre et al., 2023], which maintains a set of policies $\Pi^{\prime}\subseteq\Pi$ and iteratively chooses the weight vector $\mathbf{w}$ with the highest difference between the optimal policy value for $\mathbf{w}$ in $\Pi$ vs in $\Pi^{\prime}$. Our implementation of GPI uses the differential evolution solver from scipy for this global optimization step. The results in the table indicate that our approach achieves significantly better approximation quality. We omit other details of GPI here and refer the interested reader to [Alegre et al., 2023].
Table 3: A comparison of actual approximation ratios across various portfolio sizes for $p$-MeanPortfolio and our implementation of GPI.
Portfolio
Size
$p$-Mean-
Portfolio
GPI
Resource Allocation after Natural Disaster
1 0.706 0.514
2 0.904 0.520
3 0.999 0.520
4 0.100 0.520
Healthcare Intervention
1 0.924 0.347
2 0.982 0.641
3 0.982 0.641
4 0.982 0.641
5 0.993 0.641
6 0.999 0.641
7 1.000 0.641
Appendix D Additional Experimental Results
Figures 3 and 4 illustrate how the optimal policies for different $p$ values in the portfolio lead to varying impacts for the stakeholders. These portfolios are all obtained using $p$-MeanPortfolio. We also include the values of $p$ chosen by $p$-MeanPortfolio to construct the portfolio in Table 4.
Table 4: The set of values of $p$ chosen by $p$-MeanPortfolio to obtain the corresponding portfolios.
Portfolio
size
$p$ Values
Resource Allocation After Natural Disaster
1 -0.829
2 -4.864, -0.466
3 -11.136, -2.034, 0.621
4 -15.29, -2.054, 0.517, 0.758
Healthcare Intervention
1 -1.325
2 -2.864, 0.034
3 -4.333, -0.333, 0.333
4 -6.641, -0.433, -0.075, 0.463
5 -9.216, -4.108, -0.437, -0.078, 0.461
6 -13.801, -6.4, -0.594, -0.22, 0.085, 0.542
7 -17.793, -3.698, -0.549, -0.186, -0.038, 0.222, 0.611
Taxi Environment
1 -2.0
2 -3.89, 0.87
3 -6.21, 0.72, 0.86
8 -13.16, -10.06, -6.17, -4.82, 0.66, 0.7, 0.77, 0.89
10 -27.03, -13.02, -6.25, 0.66, 0.71, 0.78, 0.81, 0.86, 0.89, 0.95
Figure 3: Normalized total reward for each route under different policies in the portfolio generated by $p$-MeanPortfolio in the taxi environment.
Appendix E Experimental Details
@@ E.1 Taxi Environment
This environment consists of a taxi agent whose task is to deliver passengers from source to destination. The world consists of a 6x6 grid and based on the environment setup of [Fan et al., 2022], we consider 4 source-destination pairs. Whenever a taxi moves to a source point, it can pick a passenger, move to destination and drop the passenger, thus receiving a reward. Since some routes could be easier to serve, the agent has to decide which route to serve more often. Rewards are multi-dimensional, each dimension corresponding to each route. We consider the following source-destination pairs: $source\_coords=[[0,0],[0,5],[3,0],[1,0]]$, $destination\_coords=[[1,5],[5,0],[3,3],[0,3]]$
@@
The state space consists of information on the location of the taxi, location of passengers and whether passengers have been picked up by the taxi at the moment.
@@
The agent can take one of 6 actions: move north, south, east, west and pick or drop a passenger in the current grid cell.
@@
The reward function gives reward 0 for moving, -10 reward for taking invalid action of picking or dropping a passenger at wrong coordinating, and 30 reward for dropping a passenger at the right destination.
@@
Given a scalarization function (or welfare function ), the task is to find a policy $\pi^{*}$ that maximizined the expected value of scalarized return. Specifically:
@@
We learn this policy using the Welfare Q-Learning algorithm proposed in [Fan et al., 2022]. For every problem setting, we train the policy for 200 episodes. Every episode is a finite horizon problem with 1000 timesteps. We consider discount factor gamma as 0.99.
E.2 Natural Disaster
@@ 12 Low Near Middle-Income 230 100
E.3 Healthcare Intervention
The Healthcare Intervention problem is based on the large-scale mobile-health program run by the NGO ARMMAN [ARMMAN, 2024]. The goal of the program is to maximize engagement of beneficiaries with the program using limited service call interventions. Based on previous works, we model this problem as RMAB problem where we have multiple arms or beneficiaries, a budget on the number of service calls to give every week. For every beneficiary, we have information on their listenership with the voice call. This is considered as an engaging or non-engaging state if the listenership is above or below 30-seconds threshold respectively. One week of time is considered as one timestep. Finally, every week, the action can be to place (active) or not to place a live service call (passive). Additionally, for every beneficiary, we have information on their socio-demographic characteristics such as age, income, education. We use real world data from service quality improvement conducted by ARMMAN in January 2022 [Verma et al., 2023]. Further, for our experiments, we sample data for $2100$ beneficiaries, and run the experiment for $H=12$ timesteps.
Generation of Policies & Reward Functions
The default reward incentivizes agents to be in engaging state. Thus, agents receive reward 1 if they are in engaging state, and 0 otherwise. However, due to varying policy needs over time or over different geographies, health workers often have to prioritize some beneficiaries more than others. To capture these varying priorities, we leverage the Social Choice Language Model framework [Verma et al., 2024a] to derive reward functions from natural-language commands. Given a command specifying $N$preferences (each indicating a subgroup to prioritize) this method generates two sets of reward functions:
  1. 1.
    Individual preferences: one reward function per preference (total of $N$ reward functions) which we treat as each stakeholder’s reward function.
  2. 2.
    Balancing functions: a collection of reward functions that trade off among the $N$ preferences, whose corresponding policies form our feasible set $\Pi$.
In our experiments, we set $N=59$ and construct $|\Pi|=200$ balancing policies as feasible set.